July 29, 2015 Calibration
GlueX Calibration Meeting
Wednesday, July 29, 2015
11:00 am, EDT
JLab: CEBAF Center, F326
You can connect using BlueJeans using the meeting number 630 804 895 . (Click "Expand" to the right for more details -->):
- Make sure you have created a BlueJeans account via your JLab CUE account using this link:
- http://jlab.bluejeans.com (You should only need to do this once)
- Meeting ID: 630804895
- (you may need to type this in, depending how you connect)
- If connecting via Web Browser: click this link (no passcode is needed):
- If connecting via iOS or Android App:
- Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting
- If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"
- Dial: 1 888 240 2560 (US or Canada only)
- or, List of International Numbers
- If connecting via Polycom unit:
- Dial 126.96.36.199 or bjn.vc
- Enter meeting ID above
- Use *4 to unmute
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2015/ .
- Offline Monitoring Update (Kei)
- Summary of Detector Performance
- Smeared HDDM files available at /volatile/halld/detcom_02/smeared/
- REST & monitoring files being reconstructed
- Subdetector Reports
- Calibration Processing
Attending: Sean (NU); Simon, Kei, Mark I., Sebastian, Adesh, Brad, Elton, Alex B., Nathan, Lubomir (JLab); Curtis, Mike S., WIll M. (CMU)
Kei gave an update on the status of the latest run of the offline monitoring train. This was the first run to use SWIF for job submission, and the first large scale data analysis to use SWIF (as far as we know). Processing went slow this time primarily due to a high load on the cluster. Some shortcomings of SWIF were identified and Kei is in communication with Chris Larrieu (author of SWIF) to fix them. Kei is also working on wrapper for SWIF that will provide an easy-to-use interface for running over data.
See his slides for more details and statistics on this processing run.
There was some discussions about what the performance of SWIF job submission would be
Summary of Detector Performance
Curtis has only received updates from Paul, about the RF time performance. He again requested that people send him material, or update the document themselves, to have a snapshot of detector performance for this summer.
For the simulations of spring data, the smeared HDDM files are available at /volatile/halld/detcom_02/smeared/ . The REST and monitoring data are currently being recreated.
Sean pointed out that the smeared HDDM files are ready to be analyzed, for example Justin has already been using them, and requested that any feedback on them to be send to him and Mark Ito.
Adesh reported that he is updating the gains for the field off data, since previous studies had shown that the gains are different for when the magnetic field is on and off. He is currently studying the non-linear effects, and looking at simulations to help determine the proper correction.
Elton reported that George is working on finalizing his results for the effective velocities and relative time offsets between the upstream and downstream ends, and planned to report final results tomorrow. Mike S. has started to look at the TDC information in the BCAL.
Mike S. reported that no data has been taken with the CDC, since one of the mass flow controllers died and the CDC was filled with argon. Fixes are being applied to get back to the correct gas mix. With the current data he is investigating how the track residuals are affects by his straw entrance angle-dependent corrections and the z dependence of the drift time distribution effects he has been seeing.
Lubomir has been working on improving the cathode strip reconstruction, and has gotten the resolution down to 200 um. He will work with Simon to integrate these improvements into sim-recon.
Mark I. reported that Beni noticed a significant improvement in using the calibration constants in his CCDB variation compared to the current default constants. He had thought that these were the same, so he is trying to understand where this difference is coming from.
Brad Cannon, a grad student from FSU, is visiting JLab for a couple weeks, to learn the calibration procedures that Beni developed. He has been making good progress and has already found a couple of bugs, and it's expected that he will take over the calibration of the TOF.
Eric reported that he has finished his propagation time correction studies, and found that the beam data gave better corrections than the values measured on the test stand. He is working on finalizing this effort.
Mahmoud is in charge of regularizing these procedures for automation, and Sean hoped that a plan could be put together in order to move forward with that project.
Alex B. is finishing his preliminary energy calibration of the PS, using both the TAGM and TAGH.
Nathan is taking another look to see if he can improve the timing calibrations. Currently, he is looking to see if the timewalk calibrations can be improved.